# Hestia: Hierarchical Next-Best-View Exploration for Systematic Intelligent Autonomous Data Collection

### v1

@@ Abstract
Advances in 3D reconstruction and novel view synthesis have enabled efficient, photorealistic rendering, but the data collection process remains largely manual, making it time-consuming and labor-intensive. To address the challenges, this study introduces Hierarchical Next-Best-View Exploration for Systematic Intelligent Autonomous Data Collection (Hestia [^1] In Greek mythology, Hestia is the goddess of the hearth, symbolizing home, foundation, and structure, representing stability and guidance in complex systems), which leverages reinforcement learning to learn a generalizable policy for 5-DoF next-best viewpoint prediction. Unlike prior approaches, Hestia systematically defines the next-best-view task by proposing core components such as dataset choice, observation design, action space, reward calculation, and learning schemes, forming a foundation for the planner. Hestia goes beyond prior next-best-view approaches and traditional capture systems through integration and validation in a real-world setup, where a drone serves as a mobile sensor for active scene exploration. Experimental results show that Hestia performs robustly across three datasets and translated object settings in the NVIDIA IsaacLab environment, and proves feasible for real-world deployment. The code and data processing script for Hestia are included in the supplementary materials and will be released after the paper is published.
Figure 1: Data collection system. Hestia is an intelligent data collection system that integrates a next-best-view planner with a drone-based real-world platform, enabling robust viewpoint prediction even under spatial shifts of objects. Please refer to our demonstration video for further details.
Keywords: Data Collection, Machine Perception, Reinforcement Learning
@@ 1 Introduction
Throughout the years, multiview-based 3D scene reconstruction [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] and novel view synthesis [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] have been critical research fields in computer vision and graphics. These methods leverage multiview information (e.g., correspondences) from images to reconstruct high-fidelity scenes or synthesize photorealistic novel views. However, collecting multiview data is often time-consuming and labor-intensive, as these methods usually depend on manually captured data. The situation becomes more challenging for generalizable approaches [31, 23, 26, 30, 17, 29, 24] that require diverse and extensive data [31, 32, 33, 34, 35]. From a data collection perspective, one way to reduce the effort is to develop machines (e.g., multiview camera rigs) [36, 37, 38, 39, 40, 41, 42] to minimize human intervention. Although these machines make the data collection process more scalable, their sensors are usually passive (e.g., non-movable), lacking the ability to actively explore the scene. This lack of active exploration may result in missing occluded parts. To enable active capture and alleviate the effort required for manual capture, we propose Hestia, a Hierarchical Next-Best-View Exploration for Systematic Intelligent Autonomous Data Collection. The goal of Hestia is to actively collect data from object-centric scenes with a higher coverage ratio and less amount of data, demonstrating that the proposed prototype has the potential to operate in real-world settings. We address this problem by introducing two main modules, which are a next-best-view planner (see Sec. 3) that actively determines a five-degree-of-freedom (5 DoF) viewpoint (e.g., camera position, yaw, and pitch) based on voxel face observations and an intelligent data collection system (see Sec. 4) that uses a drone equipped with an RGB camera as the agent for data collection.
@@
Figure 2: A voxel is worth more than a ray. Unlike the RL-based generalizable method [43], Hestia treats each voxel as a cube by considering its six faces, rather than a point. This reduces the information loss inherent in point approximations, ensuring a more accurate representation of the voxel.
In recent decades, next-best-view planners have been developed for active capture, demonstrating promising potential for reducing the human workload. This is particularly important when a professional human pilot is unavailable to control the robot (e.g., a drone), or when human resources are limited or urgently needed elsewhere. Traditional next-best-view planners [44, 45, 46, 47, 48] rely on heuristic rules to predict the viewpoint. Although these approaches show promising and awesome performance for specific scenarios, the same handcrafted rules or hyperparameters may not be the best setting for other scenes as discussed by [43]. Thanks to advancements in deep learning and computational capabilities, several learning-based (e.g., online learning or generalizable methods) next-best-view planners [49, 50, 51, 52, 53, 54, 43, 55, 56, 57, 58] have been developed. Among these approaches, a generalizable RL-based method [43] that utilizes an occupancy grid as the observation demonstrates effective performance in terms of coverage ratio, viewpoint flexibility, and generalization capability.
@@
Inspired by the approach [43], we develop a next-best-view planner utilizing a voxel grid as the observation. Unlike the prior work [43], our method introduces a concept, ”A voxel is worth more than a ray”, by incorporating the visibility of the six faces of each voxel into the observation and reward function (see Fig. 2 and Sec. E). Theoretically, if we keep sampling with a 1-ray camera in a scene containing $k$ unit cubes, then from a coupon collector’s perspective [59], treating each voxel as a point will result in approximately $k^{-\frac{1}{6}}$ of the faces being missed when sampling stops (see appendix Sec. A). In contrast, treating each voxel as a cube allows faces to be fully covered. Hence, Hestia accounts for the visibility of individual voxel faces and achieves more comprehensive data capture. In addition, representing each voxel as a cube does not introduce significant computational overhead and still enables real‑time operation at 25 FPS (see Tab. 1).
@@
To further help the learning process, Hestia focuses on three key aspects of generalizable RL: the observation (e.g., the diversity and scale of the dataset), the action (e.g., the dimension of the action space), and the learning process (e.g., strategies for training). For observation, Hestia utilizes the largest dataset processed from Objaverse [60, 61], for the next-best-view task (see Sec. I). Therefore, Hestia encounters a more diverse range of surface geometries, rather than just cubic shapes (see Sec. I). For action, instead of directly predicting the 5 DoF next-best viewpoint, a task that is inherently complex and challenging, Hestia adopts a hierarchical structure to simplify the process. The approach begins by predicting the look-at point, which serves as the target of focus, followed by determining the viewpoint position based on this point. Finally, Hestia formulates the next-best-view task as a greedy optimization problem. Given an occupancy grid, the goal is to find the viewpoint that maximizes the coverage ratio. To achieve this, Hestia relies solely on the previous image, camera pose, and the given occupancy grid to predict the next-best-view, rather than depending on a sequence of images [55, 43] and camera poses [43]. Additionally, the reward discount factor $\gamma$ is set to a small value to prioritize current improvements without relying on a disproportionately large goal reward. Following a greedy algorithm, Hestia can mitigate the spurious correlations [^2] Spurious correlations [62, 63] refer to certain groups contributing to model errors. In this study, spurious correlations refer to large positive future rewards assigned to suboptimal current next-best-view decisions, leading to ineffective policy learning. (see Fig. A7) arising from false associations between the current action and a large future reward (e.g., the final goal reward). Hence, Hestia can reach a better coverage ratio with fewer images than previous methods [53, 43, 55].
The proposed next-best-view planner is integrated into an intelligent data collection system, with the planner running on a ground station, to demonstrate its effectiveness in real-world scenarios. The system employs a drone as an agent, equipped with a camera mounted on a gimbal, to capture data. Using drones instead of robot arms [51] practically enables data collection in regions that are difficult for humans to access. The contributions of this work are:
@@
  • A generalizable RL-based next-best-view planner achieves a better coverage ratio with fewer images by explicitly considering voxels as cubes rather than points.
  • Propose a hierarchical structure to handle the high-dimensional action space, use the largest dataset to ensure robustness, and adopt a greedy algorithm to reduce spurious correlations.
  • Integrate the next-best-view planner into an intelligent, autonomous data collection prototype to demonstrate its potential for active data capture under real-world scenarios.
@@ 2 Related Work
Data collection systems. Advances in 3D reconstruction [64, 8, 9, 10, 11, 12, 13, 14, 15, 65, 66] and novel view synthesis [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 67, 68, 69, 70, 71, 72] techniques have enabled the rendering of high-fidelity and photo-realistic scenes. However, most approaches depended on manually captured data [73, 39, 74, 75], which is both time-consuming and labor-intensive. This limitation makes the data collection process less practical for generalizable approaches [31, 23, 26, 30, 17, 29, 24, 70, 71, 13, 14, 72, 66]. To overcome this, some methods [60, 61, 31, 35, 34, 76, 77] collected data from the internet, leveraging contributions from people worldwide. While these approaches enable the rapid collection of large datasets and benefit generalizable methods, they face potential drawbacks such as inconsistent data quality and the risk of saturation due to the inability to generate their own data. Instead, many people developed their data collection machines [36, 37, 38, 39, 40, 41, 78, 79, 80] to capture high-quality data, which is a more sustainable approach in the long term. For instance, multiview camera rigs have been used to capture face-forward scenes [37, 38, 39, 40, 80], 360° outward-facing scenes [36], and 360° inward-facing scenes [41]. Other studies [79] have combined multiple camera devices (e.g., spectacles camera and 360° camera) to capture scenes from different perspectives. Additionally, ego-centric approaches [78] developed wearable camera glasses to collect data from a first-person perspective while the user was in motion. Although these methods can collect large amounts of data using multiple cameras or passively gather data through human-operated devices, they lack the capability to actively explore the scene. In contrast, our proposed data collection system is intelligent, enabling autonomous exploration of the scene using a drone as a mobile sensor.
@@
Scene-specific next-best-view planners. Next-best-view planners have demonstrated promising results in active 3D reconstruction by predicting the optimal viewpoint for data capture based on the current state. Traditional approaches [44, 45, 46, 47, 48] rely on hand-crafted rules to determine the next-best viewpoint. For instance, the method [45] selected the next-best viewpoint by maximizing a rating function favoring smooth regions, which may overlook fine-grained object details. Instead, the methods [44, 47] collected data along boundaries between seen and unseen surfaces to capture finer details, but still require handcrafted parameter tuning for each scene. Another approach [46] scanned segmented objects sequentially using a predefined object database, reducing the need for handcrafted tuning, but its performance degrades in cluttered environments due to inaccurate object matching. Recent advances in deep learning and increased computational power have given rise to learning-based methods [49, 50, 51, 52, 53, 54, 43, 55, 56]. Some studies [49, 54] used NeRF [16] ensembles to estimate uncertainty via model disagreement for viewpoint selection, resulting in linearly increasing computational overhead. Other works [50, 56] avoid the computational overhead by incorporating Bayesian-based NeRF [81, 82] to estimate uncertainty for viewpoint selection. Meanwhile, other approaches [51, 52] defined the next-best viewpoint as the viewpoint that maximizes the entropy of the density field along the camera rays. Although these methods have shown outstanding performance in collecting data, all require sampling candidate viewpoints. In addition, their reliance on online learning makes them less suitable for real-time applications.
@@
Generalizable next-best-view planners. Unlike the aforementioned online-learning approaches, the generalizable methods [53, 43, 55] avoid the training process for new scenes, thereby enabling faster next-best-view selection. The prediction time is important for real-world tasks where a robot (e.g., a drone) may run out of battery within a few minutes. Among generalizable methods, prior work [53] proposed a Bayesian-based NeRF that selects the next-best viewpoint to maximize view variance without additional training. However, it still requires candidate viewpoint sampling, resulting in non-real-time performance. Instead, another line of generalizable next-best-view approaches [43, 55] utilized reinforcement learning (RL) to learn a next-best-view planner to bypass the need for sampling candidate viewpoints. Prior work [55] proposed learning a 3-DoF next-best-view policy using a series of grayscale images as observations. Subsequently, prior work [43] improved upon this method by incorporating occupancy grids into the observations, which provide explicit geometric information. This enhancement enabled the development of a 5-DoF next-best-view planner, achieving an outstanding coverage ratio for unknown scenes.
@@
As shown in Tab. A4, compared to scene-specific methods, Hestia does not require candidate viewpoint sampling or inference-time optimization, thereby enabling more flexible viewpoint prediction. Compared to the generalizable method [43], Hestia treats voxels as cubes rather than points, enabling more comprehensive capture. In addition, Hestia adopts a greedy scheme to mitigate spurious correlations, introduces a hierarchical structure to model the action space, and uses a more diverse dataset to maintain robust performance across varying object shapes and positions.
@@ 3.1 RL Problem Definition
Our next-best-view task is to identify a 5-DoF viewpoint that maximizes the coverage ratio of an incomplete occupancy grid of the scene. The task’s goal is similar to greedy methods, which always seek the locally optimal solution. We formulate the problem as a Markov Decision Process (MDP), denoted by the tuple $\{S,A,P,R,\gamma\}$. At each time step $t$, the agent with an RGB-D camera observes a state $s_{t}$ from the set of all possible states $S$ and chooses an action $a_{t}$ from the action space $A$. The environment then transitions to the subsequent state $s_{t+1}$ according to the probabilities described by $P$, and provides a reward $r_{t}$. The magnitude of this reward is determined by the reward function $R(\cdot\mid s,a):S\times A\to r$. In reinforcement learning (RL), the main goal is to discover an optimal policy $\pi$ that maximizes the expected sum of discounted rewards, given by $E_{t}=\sum_{k=0}^{\infty}\gamma^{k}r_{t+k},$ where $\gamma\in(0,1)$ is the discount factor. We set $\gamma$ to 0.1 to align with the greedy-like objective and to avoid spurious correlations from large positive future rewards (see Fig. A7).
@@
State space. The state space of Hestia is defined as $S=\Big{\{}s_{t}\;\Big{|}\;s_{t}=\big{\{}I_{t},M_{t},G_{t},L_{t}\big{\}},\;t\in\mathbb{N}\Big{\}},$ where $I_{t}\in\mathbb{R}^{h\times w}$ is the grayscale image with height $h$ and width $w$, and $L_{t}\in\mathbb{R}^{3}$ is the camera look-at point. The vector $M_{t}\in\mathbb{R}^{6}$ consists of $X_{t}\in\mathbb{R}^{5}$, which is the camera position, pitch, and yaw, as well as $H_{t}\in\mathbb{R}^{1}$, representing the maximum flyable height for the capture. Meanwhile, $G_{t}\in\mathbb{R}^{g\times g\times g\times 10}$ includes the aggregated grid information at resolution $g$, consisting of $O_{t}\in\mathbb{R}^{g\times g\times g\times 1}$ for the cumulative occupancy grid, $C_{t}\in\mathbb{R}^{g\times g\times g\times 3}$ for the positional encoding, and $F_{t}\in\{0,1\}^{g\times g\times g\times 6}$ for the cumulative face visibility. The cumulative face visibility is updated iteratively as $F_{t}=f_{t}\lor F_{t-1},$ where $F_{t}$ represents the cumulative face visibility for all voxels up to time $t$, and $f_{t}\in\{0,1\}^{g\times g\times g\times 6}$ denotes the current face visibility. To compute $f_{t}$, the depth image $D_{t}$ is unprojected into a voxelized point cloud $V=\{v_{i}\;|\;i\in\mathbb{N}\}$, where $v_{i}$ is the $i$-th voxel. Each voxel $v_{i}$ is associated with a viewing direction vector $d_{v_{i}}\in\mathbb{R}^{3}$, defined as the vector pointing from the voxel center to the collision-free camera position $a^{{}^{\prime}}_{t}$. The vector $\mathbf{d}_{v_{i}}$ is computed as $\mathbf{d}_{v_{i}}=\frac{a^{{}^{\prime}}_{t}-p_{v_{i}}}{\|a^{{}^{\prime}}_{t}-p_{v_{i}}\|},$ where $p_{v_{i}}$ is the center of voxel $v_{i}$. For each voxel $v_{i}$ and its six outward-facing face normals $n_{i,j}\in\mathbb{R}^{3}$, the face visibility is determined and aggregated as
$f_{t}(v_{i},j)=\mathbbm{1}\left(d_{v_{i}}\cdot n_{i,j}>0\right),\quad\forall v_{i}\in V,\,j\in\{1,\dots,6\}$ (1)
@@
where $\mathbbm{1}(\cdot)$ is the indicator function. By iterating over all voxels and their respective faces, $f_{t}$ is constructed, and the cumulative visibility $F_{t}$ is updated accordingly. Although this method cannot handle all face visibilities, the approximation enables efficient computation of face visibility. Unlike prior works [43], which consider only $O_{t}$ and $C_{t}$ and thereby treat voxels as points, we treat voxels as cubes to mitigate the information loss caused by approximating voxels as points (see Fig. 2). For details regarding $O_{t}$ and $C_{t}$, please refer to the work [43].
@@
Action space. The action space $A=\Bigl{\{}a_{t}\;\Big{|}\;a_{t}\in[-1,1]^{3},\;t\in\mathbb{N}\Bigr{\}}$ represents the set of possible 3-DoF viewpoints (e.g., camera positions) at each time step $t$, where each coordinate is initially bounded within [1,1][-1,1][ - 1 , 1 ]. These coordinates are subsequently normalized to the environment’s scale to ensure appropriate positioning within the scene. Additionally, the camera’s pitch and yaw are derived from the look-at point and the collision-free action $a^{{}^{\prime}}_{t}$ converted from $a_{t}$ (see Sec. 3.2).
@@
Reward. The reward function is defined as $r_{t}=R(s_{t},a_{t})=r_{\text{coverage}}(s_{t},a_{t})+r_{\text{constraint}}(s_{t},a_{t}),$ where $r_{\text{coverage}}(s_{t},a_{t})$ encourages the observation of new voxel faces and is expressed as
$r_{\text{coverage}}(s_{t},a_{t})=\frac{\sum_{i=1}^{N}\sum_{j=1}^{6}\left(F_{t}^{i,j}-F_{t-1}^{i,j}\right)\cdot M_{\text{col}}}{N\cdot 6}\cdot 0.3$ (2)
@@
where $F_{t}^{i,j}$ and $F_{t-1}^{i,j}$ represent the visibility status of the $j$-th face of the $i$-th voxel at time $t$ and $t-1$, respectively. Here, $M_{\text{col}}\in\{0,1\}$ is a collision indicator, set to 0 in the event of a collision, thereby preventing any positive reward for invalid actions. The term $r_{\text{constraint}}(s_{t},a_{t})=-0.01$ is applied when unsafe or invalid actions occur (see Sec. E for details). Our reward is based on the face coverage ratio rather than the point coverage ratio to ensure more comprehensive capture (see Fig. 2). Furthermore, to prevent spurious correlations, the reward design aligns with a greedy-like objective, which differs significantly from prior works [43, 55] that provide a large goal reward when the coverage ratio reaches a predefined target.

### v2

# Hestia: Voxel-Face-Aware Hierarchical Next-Best-View Acquisition for Efficient 3D Reconstruction

@@ Abstract
Advances in 3D reconstruction and novel view synthesis have enabled efficient and photorealistic rendering. However, images for reconstruction are still either largely manual or constrained by simple preplanned trajectories. To address this issue, recent works propose generalizable next-best-view planners that do not require online learning. Nevertheless, robustness and performance remain limited across various shapes. Hence, this study introduces Voxel-Face-Aware Hierarchical Next-Best-View Acquisition for Efficient 3D Reconstruction (Hestia [^1] In Greek mythology, Hestia is the goddess of the hearth, symbolizing home, foundation, and structure, representing stability and guidance in complex systems), which addresses the shortcomings of the reinforcement learning-based generalizable approaches for five-degree-of-freedom viewpoint prediction. Hestia systematically improves the planners through four components: a more diverse dataset to promote robustness, a hierarchical structure to manage the high-dimensional continuous action search space, a close-greedy strategy to mitigate spurious correlations, and a face-aware design to avoid overlooking geometry. Experimental results show that Hestia achieves non-marginal improvements, with at least a 4% gain in coverage ratio, while reducing Chamfer Distance by 50% and maintaining real-time inference. In addition, Hestia outperforms prior methods by at least 12% in coverage ratio with a 5-image budget and remains robust to object placement variations. Finally, we demonstrate that Hestia, as a next-best-view planner, is feasible for the real-world application. Our project page is https://johnnylu305.github.io/hestia_web.
@@ 1 Introduction
Multiview-based 3D scene reconstruction [schoenberger2016sfm, yao2018mvsnet, xie2019pix2vox, murez2020atlas, wang2021multi, sayed2022simplerecon, xu2023unifying, wang2024dust3r, duisterhof2024mast3r, pan2024glomap] and novel view synthesis [mildenhall2020nerf, yu2021pixelnerf, yu2022plenoxels, mueller2022instant, zheng2024gpsgaussian, gslrm2024, charatan2024pixelsplat, sabour2024spotlesssplats, flynn2024quark, chen2025mvsplat] have been central topics in computer vision. These methods leverage multiview information from images to reconstruct high-fidelity scenes or synthesize photorealistic novel views. However, data acquisition remains a bottleneck. Most data is collected manually, which is time-consuming and labor-intensive, or follows preplanned camera trajectories that do not generalize well to diverse objects.
To reduce human effort, next-best-view (NBV) planning has been explored for active capture [monica2018contour, zha1997next, liu2018object, hardouin2020surface, hardouin2020next, lin2022active, pan2022activenerf, lee2022uncertainty, zhan2022activermap, jin2023neu, sunderhauf2023density, ran2023neurar, guedon2023macarons, jiang2023fisherrf]. Traditional next-best-view methods rely on heuristic rules that can work well in specific scenarios but often fail to transfer because fixed rules or hyperparameters do not adapt across scenes [chen2024gennbv]. Learning-based next-best-view planners, including online-learning and generalizable methods, improve over preplanned trajectories, which frequently miss occluded regions. Within learning-based approaches, reinforcement learning-based (RL-based) generalizable methods [peralta2020next, chen2024gennbv], which pretrain on a dataset to avoid online learning and directly predict viewpoints as actions, show promising results. This removes candidate-viewpoint sampling, which may potentially miss the best views and slow viewpoint acquisition. An occupancy-grid formulation [chen2024gennbv] further demonstrates strong coverage, viewpoint flexibility, and generalization. Nevertheless, performance remains limited and insufficiently robust across diverse object geometries.
To address the shortcomings, we propose Hestia, a Voxel-Face-Aware Hierarchical Next-Best-View Acquisition for Efficient 3D Reconstruction. Hestia actively collects data in object-centric scenes by predicting five-degree-of-freedom (5-DoF) viewpoints (x, y, z, yaw, pitch) from voxel-face observations. Specifically, Hestia systematically defines the next-best-view task by proposing core components such as dataset choice, observation and reward design, action space, and learning schemes, forming a foundation for the planner.
@@
Figure 1: A voxel is worth more than a ray. Unlike the RL-based generalizable method [chen2024gennbv], Hestia treats each voxel as a cube by considering its six faces, rather than a point. This reduces the information loss inherent in point approximations, ensuring a more accurate representation of the voxel.
@@
An idea is “A voxel is worth more than a ray”. We incorporate the visibility of the six faces of each voxel into both the observation and the reward function (see Figs. 1 and 3). Theoretically, if we sample with a one-ray camera in a scene with $k$ unit cubes and treat each voxel as a point, then from a coupon collector’s perspective [boneh1997coupon] approximately $k^{-1/6}$ of the faces will be missed when sampling stops (see Sec. S1). Treating each voxel as a cube enables full face coverage, so Hestia accounts for individual voxel-face visibility and captures data more comprehensively. This representation adds little computational overhead and still achieves real-time operation at 25 FPS (see Tab. 1).
@@
Hestia further improves learning by refining the observation, action, and learning process. For observation, Hestia uses the largest dataset that we processed from Objaverse [Deitke_2023_CVPR, deitke2024objaverse] for the next-best-view task, exposing the policy to a broad range of surface geometries rather than mostly cubic shapes (see Sec. S8). For action, instead of predicting the full 5-DoF next-best view in one step, Hestia adopts a hierarchical structure. The policy first predicts a look-at point as the target of attention, then determines the viewpoint position conditioned on this point. For learning, Hestia formulates the task as a close-greedy optimization problem in which, given an occupancy grid, it selects the view that maximizes the current coverage ratio. The policy relies only on the previous image, the previous camera pose, and the occupancy grid, rather than a long sequence of images and poses [peralta2020next, chen2024gennbv]. We also use a small reward discount factor $\gamma$ to prioritize immediate improvements without depending on an oversized terminal reward. Similar to a greedy algorithm, this reduces spurious correlations [^2] Spurious correlations [interpretation1971spurious, kim2024discovering] refer to certain groups contributing to model errors. In this study, spurious correlations refer to large positive future rewards assigned to suboptimal current next-best-view decisions, leading to ineffective policy learning. between current actions and large future rewards (see Fig. S8). As a result, Hestia reaches higher coverage with fewer images than prior methods [guedon2023macarons, jin2023neu, chen2024gennbv, peralta2020next]. Finally, we demonstrate that Hestia, as a next-best-view planner, is feasible for real-world application using a drone with an RGB camera as a mobile agent and a depth predictor [duisterhof2024mast3r, wang2024dust3r] to convert RGB images into depth maps. The contributions of this work are as follows:
@@
  • A RL-based generalizable next-best-view planner that considers voxels as cubes rather than points to avoid geometry overlooking.
  • A hierarchical structure for handling the high-dimensional continuous action space, a larger and more diverse training set for promoting robustness, and a close-greedy strategy for reducing spurious correlations.
  • Comprehensive evaluations on three datasets show that Hestia achieves non-marginal improvements and is suitable for 3D reconstruction under limited acquisition budgets.
@@ 2 Related Work
The literature review mainly focuses on next-best-view methods that are formulated in 5 DoF or assume a drone as an agent.
@@
Scene-specific next-best-view planners. Next-best-view planners have demonstrated promising results in active 3D reconstruction by predicting the optimal viewpoint for data capture based on the current state. Traditional approaches [monica2018contour, zha1997next, liu2018object, hardouin2020surface, hardouin2020next, dai2020fast, zhou2021fuel] rely on hand-crafted rules to determine the next-best viewpoint. For instance, the method [zha1997next] selected the next-best viewpoint by maximizing a rating function favoring smooth regions, which may overlook fine-grained object details. Instead, the methods [monica2018contour, hardouin2020surface] collected data along boundaries between seen and unseen surfaces to capture finer details, but still require handcrafted parameter tuning for each scene. Another approach [liu2018object] scanned segmented objects sequentially using a predefined object database, reducing the need for handcrafted tuning, but its performance degrades in cluttered environments due to inaccurate object matching. Recent advances in deep learning and increased computational power have given rise to learning-based methods [lin2022active, pan2022activenerf, lee2022uncertainty, zhan2022activermap, jin2023neu, sunderhauf2023density, chen2024gennbv, peralta2020next, ran2023neurar, zeng2022efficient, tao2022seer]. Some studies [lin2022active, sunderhauf2023density] used NeRF [mildenhall2020nerf] ensembles to estimate uncertainty via model disagreement for viewpoint selection, resulting in linearly increasing computational overhead. Other works [pan2022activenerf, ran2023neurar, zeng2022efficient] avoid the computational overhead by incorporating Bayesian-based NeRF [shen2021stochastic, martin2021nerf] to estimate uncertainty for viewpoint selection. Meanwhile, other approaches [lee2022uncertainty, zhan2022activermap] defined the next-best viewpoint as the viewpoint that maximizes the entropy of the density field along the camera rays. Although these methods have shown outstanding performance in collecting data, all require sampling candidate viewpoints. In addition, their reliance on online learning makes them less suitable for real-time applications.
@@
Generalizable next-best-view planners. Unlike the aforementioned online-learning approaches, the generalizable methods [jin2023neu, chen2024gennbv, peralta2020next] avoid the training process for new scenes, thereby enabling faster next-best-view selection. Prediction time is important for real-world tasks where a robot (e.g., a drone) may run out of battery within a few minutes. Among generalizable methods, prior work [jin2023neu] proposed a Bayesian-based NeRF that selects the next-best viewpoint to maximize view variance without additional training. However, it still requires candidate viewpoint sampling, resulting in performance unsuitable for real-time applications. Instead, another line of generalizable next-best-view approaches [chen2024gennbv, peralta2020next] utilized reinforcement learning to learn a next-best-view planner to bypass the need for sampling candidate viewpoints. Prior work [peralta2020next] proposed learning a 3-DoF next-best-view policy using a series of grayscale images as observations. Subsequently, prior work [chen2024gennbv] improved upon this method by incorporating occupancy grids into the observations, which provide explicit geometric information. This enhancement enabled the development of a 5-DoF next-best-view planner, achieving an outstanding coverage ratio for unknown scenes.
@@
As shown in Tab. S4, compared to the scene-specific methods, Hestia does not require candidate viewpoint sampling or inference-time optimization, thereby enabling more flexible viewpoint prediction. Compared to the generalizable method [chen2024gennbv], Hestia treats voxels as cubes rather than points, enabling more comprehensive capture. In addition, Hestia adopts a close-greedy scheme to mitigate spurious correlations, introduces a hierarchical structure to model the action space, and uses a more diverse dataset to maintain robust performance across varying object shapes and positions. Notably, Hestia’s hierarchical structure addresses the challenge of high-dimensional continuous action search in RL-based generalizable next-best-view planning. This differs from traditional methods [monica2018contour, hardouin2020surface, dai2020fast, zhou2021fuel], which use hierarchical structures to traverse frontiers or optimize tour efficiency.
@@ 3.1 RL Problem Definition
Our next-best-view task is to identify a 5-DoF viewpoint that maximizes the coverage ratio of an incomplete occupancy grid of the scene. The task’s goal is similar to greedy methods, which always seek the locally optimal solution. We formulate the problem as a Markov Decision Process (MDP), denoted by the tuple $\{S,A,P,R,\gamma\}$. At each time step $t$, the agent with an RGB-D camera observes a state $s_{t}$ from the set of all possible states $S$ and chooses an action $a_{t}$ from the action space $A$. The environment then transitions to the subsequent state $s_{t+1}$ according to the probabilities described by $P$, and provides a reward $r_{t}$. The magnitude of this reward is determined by the reward function:
$R(\cdot\mid s,a):S\times A\to r.$ (1)
In reinforcement learning, the main goal is to discover an optimal policy $\pi$ that maximizes the expected sum of discounted rewards, given by:
$E_{t}=\sum_{k=0}^{\infty}\gamma^{k}r_{t+k},$ (2)
where $\gamma\in(0,1)$ is the discount factor. We set $\gamma$ to 0.1 to align with the greedy-like objective and to avoid spurious correlations from large positive future rewards (see Fig. S8).
State space. The state space of Hestia is defined as:
$S=\Big\{s_{t}\;\Big|\;s_{t}=\big\{I_{t},M_{t},G_{t},L_{t}\big\},\;t\in\mathbb{N}\Big\}$ (3)
@@
where $I_{t}\in\mathbb{R}^{h\times w}$ is the grayscale image with height $h$ and width $w$, and $L_{t}\in\mathbb{R}^{3}$ is the camera look-at point. The vector $M_{t}\in\mathbb{R}^{6}$ consists of $X_{t}\in\mathbb{R}^{5}$, which is the camera position, pitch, and yaw, as well as $H_{t}\in\mathbb{R}^{1}$, representing the maximum flyable height for the capture. Meanwhile, $G_{t}\in\mathbb{R}^{g\times g\times g\times 10}$ includes the aggregated grid information at resolution $g$, consisting of $O_{t}\in\mathbb{R}^{g\times g\times g\times 1}$ for the cumulative occupancy grid, $C_{t}\in\mathbb{R}^{g\times g\times g\times 3}$ for the positional encoding, and $F_{t}\in\{0,1\}^{g\times g\times g\times 6}$ for the cumulative face visibility. The cumulative face visibility is updated iteratively as:
$F_{t}=f_{t}\lor F_{t-1}$ (4)
where $F_{t}$ represents the cumulative face visibility for all voxels up to time $t$, and $f_{t}\in\{0,1\}^{g\times g\times g\times 6}$ denotes the current face visibility. To compute $f_{t}$, the depth image $D_{t}$ is unprojected into a voxelized point cloud $V=\{v_{i}\;|\;i\in\mathbb{N}\}$, where $v_{i}$ is the $i$-th voxel. Each voxel $v_{i}$ is associated with a viewing direction vector $d_{v_{i}}\in\mathbb{R}^{3}$, defined as the vector pointing from the voxel center to the collision-free camera position $a^{{}^{\prime}}_{t}$. The vector $\mathbf{d}_{v_{i}}$ is computed as:
$\mathbf{d}_{v_{i}}=\frac{a^{{}^{\prime}}_{t}-p_{v_{i}}}{\|a^{{}^{\prime}}_{t}-p_{v_{i}}\|}$ (5)
where $p_{v_{i}}$ is the center of voxel $v_{i}$. For each voxel $v_{i}$ and its six outward-facing face normals $n_{i,j}\in\mathbb{R}^{3}$, the face visibility is determined and aggregated as
$f_{t}(v_{i},j)$ $=\mathbbm{1}\!\left(d_{v_{i}}\cdot n_{i,j}>0\right),$
$\quad\forall v_{i}\in V,\;j\in\{1,\dots,6\}$ (6)
@@
where $\mathbbm{1}(\cdot)$ is the indicator function. By iterating over all voxels and their respective faces, $f_{t}$ is constructed, and the cumulative visibility $F_{t}$ is updated accordingly. Although this method cannot handle all face visibilities, the approximation enables efficient computation of face visibility. Unlike prior works [chen2024gennbv], which consider only $O_{t}$ and $C_{t}$ and thereby treat voxels as points, we treat voxels as cubes to mitigate the information loss caused by approximating voxels as points (see Fig. 1). For details regarding $O_{t}$ and $C_{t}$, please refer to the work [chen2024gennbv].
Action space. The action space:
$A=\Bigl\{a_{t}\;\Big|\;a_{t}\in[-1,1]^{3},\;t\in\mathbb{N}\Bigr\}$ (7)
@@
represents the set of possible 3-DoF viewpoints (e.g., camera positions) at each time step $t$, where each coordinate is initially bounded within [1,1][-1,1]. These coordinates are subsequently normalized to the environment’s scale to ensure appropriate positioning within the scene. Additionally, the camera’s pitch and yaw are derived from the look-at point and the collision-free action $a^{{}^{\prime}}_{t}$ converted from $a_{t}$ (see Sec. 3.2).
Reward. The reward function is defined as:
$r_{t}=R(s_{t},a_{t})=r_{\text{coverage}}(s_{t},a_{t})+r_{\text{constraint}}(s_{t},a_{t})$ (8)
@@
where $r_{\text{coverage}}(s_{t},a_{t})$ encourages the observation of new voxel faces and is expressed as:
$r_{\text{coverage}}(s_{t},a_{t})$ $=\frac{\sum_{i=1}^{N}\sum_{j=1}^{6}\left(F_{t}^{i,j}-F_{t-1}^{i,j}\right)\cdot M_{\text{col}}}{N\cdot 6}$
$\quad\cdot 0.3$ (9)
@@
where $F_{t}^{i,j}$ and $F_{t-1}^{i,j}$ represent the visibility status of the $j$-th face of the $i$-th voxel at time $t$ and t1t-1, respectively. Here, $M_{\text{col}}\in\{0,1\}$ is a collision indicator, set to 0 in the event of a collision, thereby preventing any positive reward for invalid actions. The term $r_{\text{constraint}}(s_{t},a_{t})=-0.01$ is applied when unsafe or invalid actions occur (see Sec. S5 for details). Our reward is based on the face coverage ratio rather than the point coverage ratio to ensure more comprehensive capture (see Fig. 1). Furthermore, to prevent spurious correlations, the reward design aligns with a greedy-like objective, which differs significantly from prior works [chen2024gennbv, peralta2020next] that provide a large goal reward when the coverage ratio reaches a predefined target.